general physics

A penny is dropped form the observation deck of the empire state building (369m above the ground). With what velocity does it strike the ground?
Answer #1
vf^2 = vi^2 + 2ad

vf^2 = 0^2 + 2*-9.81*-369

vf= -85.08 m/s
answered by: Vidwatta
Answer #2

vf = √(2gd) = √(2 * 369 * 9.8) = 85m/s

answered by: Jacod
Answer #3

v*v = 2*9.8*369

v = 85.04 m/sec
answered by: Lib